Materials $6.50
Direct labor [all variable] 5.80
Total overhead 5.30
Total $17.60


$78,000 of X Company's total overhead costs are fixed; $50,700 of the $78,000 are unavoidable even if it buys the part. Also, if X Company buys the part, it can rent the equipment that was used to make the part to another company for $80,000.

3. If X Company buys the part instead of making it, it will save

Tries 0/3

4. At what production level would X company be indifferent between making and buying the part? Questions 3 and 4 refer to the following information: X Company is considering buying a part supply this part for $18.03 per
Answer #1

3) Differential analysis

Make Buy
Direct material 338000
Direct labor 301600
Variable overhead 197600
Fixed overhead 27300
Opportunity Cost 80000
Purchase cost 937560
Total relevant cost 944500 937560

it will save = 944500-937560 = 6940

Indifferent point

Manufacturing cost = Purchase cost

16.10X+107300 = 18.03X

-1.93X = -107300

X(Unit) = 55596 Units